OpenStudy (ccchristian1):

A pencil at a stationery store costs $1, and a pen costs $1.50. Shawn spent $12 at the store. He bought a total of 9 items. Which system of equations can be used to find the number of pencils (x) and pens (y) he bought? 1.5x + y = 12 x = 9y x + 1.5y = 12 x + y = 9 x + 9y = 12 x = 1.5y 9x + y = 12 x = 1.5y

OpenStudy (ccchristian1):

@hysusonic

OpenStudy (hysusonic):

so you're supposed to have 1x + 1.50y =12

OpenStudy (hysusonic):

that's how much he spent and how many items he bought would be x+y=9

OpenStudy (hysusonic):

so which one looks like that?

OpenStudy (ccchristian1):

A?

OpenStudy (anonymous):

u can either use the 1st and 2nd one...or the 3rd and 1st one

OpenStudy (hysusonic):

no... A is saying that a pencil costs $1.50 when a pencil costs $1

OpenStudy (ccchristian1):

1.5x + y = 12 x = 9y x + 1.5y = 12 x + y = 9 x + 9y = 12 x = 1.5y 9x + y = 12 x = 1.5y

OpenStudy (ccchristian1):

so which one ?

OpenStudy (hysusonic):

well look at the formulas... 1x+1.5y=12 x+y=9

OpenStudy (hysusonic):

which one looks like that?

OpenStudy (ccchristian1):

B?

OpenStudy (hysusonic):

yup

OpenStudy (anonymous):

i quess b...will do

OpenStudy (anonymous):

rite

OpenStudy (anonymous):

#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar#sayNOtoPERFUMESotherTHANattar